# The Galactic population of canonical pulsars II

### v1

@@ 2.2 Pulsar evolution
The pulsar initial period, inclination angle and magnetic field must be evolved in time self-consistently due to electromagnetic radiation and internal dissipation within the neutron star. To this end we employ the pulsar force-free magnetosphere model taking into account the plasma current and charge within the magnetosphere. The spin down luminosity $\dot{E}$ is
@@
where $I$ is the neutron star moment of inertia its value being approximately $I\approx 10^{38}$ kg.m2, $\Omega=2\pi/P$ is the rotation frequency of the pulsar ($P$ is the spin period) and $\dot{\Omega}$ is its time derivative, $R=12$ km is the typical radius of a neutron star as found by recent NICER observations (Riley et al., 2019; Bogdanov et al., 2019), $B$ is the magnetic field at the magnetic equator, $\alpha$ is the inclination angle, $\mu_{0}$ is the vacuum permeability constant, $\mu_{0}=4\pi\times 10^{-7}$ H/m and $c$ is the speed of light. The term $L_{\rm ffe}$ for the spin down luminosity was given by Spitkovsky (2006) and Pétri (2012). Equation (5) clearly shows the correlation between the obliquity, the magnetic field and the rotation frequency. Combining equation (5) and (6) leads to
@@
where quantities with index 0 indicate their initial value. The quantities without the index are their current time value.
@@
where $\alpha_{d}$ is a constant parameter controlling the speed of the magnetic field decay and $\tau_{d}=k_{\tau_{d}}\tau_{v}$ is the magnetic field decay time scale. The decay rate $\tau_{v}$ was chosen randomly between 3 values : $5\times 10^{4}$ yr or $7\times 10^{4}$ yr or $5\times 10^{5}$ yr, with a probability of 0.23, 0.46 and 0.31 respectively (these probabilities were chosen as best educated guesses after several run with different values) for each pulsar, corresponding to the $\tau_{v}$ value for canonical pulsars in Viganò et al. (2013). The values of $\alpha_{d}$ and $k_{\tau_{d}}$ can be found in Table 5 of Sect. 3.
In line with the spin down luminosity, the inclination angle satisfies another evolution equation that after integration was found by Phillipov et al. (2014) for a spherically symmetric neutron star with a constant magnetic field. With our decaying prescription the inclination angle $\alpha$ is found by solving for the root of
@@
where $t$ is the time representing the age of the pulsar and $K={R^{6}}/{I\,c^{3}}$. The typical decay timescale for a mean magnetic field of $2.5\times 10^{8}$ T is $4.6\times 10^{5}$ yr in Viganò (2013), however in this study the typical decay timescale associated to this magnetic field is $3.5\times 10^{5}$ yr which is consistent with the estimate of Viganò (2013) for the same field strength.
Finally, we need to follow the particle motion within the Galactic potential. Each generated pulsar evolves in the gravitational potential $\Phi$ subject to an acceleration ${\ddot{x}}$ according to
@@
Equation (12) is integrated numerically thanks to a Position Extended Forest Ruth-Like (PEFRL) algorithm (Peter Young, 2014), a fourth order integration scheme, presented in Appendix A with a convergence and accuracy study. The galactic potential model used in discussed in the next section.
2.3 Description of the Galactic potential
@@
The potential for the dark matter halo $\Phi_{h}$, according to the frequently used potential of Navarro et al. (1997), it is expressed as
@@ 2.4 The death line
Several equations for the death line have been proposed in the literature (Chen & Ruderman, 1993; Zhang et al., 2000; Gil & Mitra, 2001; Mitra et al., 2019). The most suitable one in our study is that of Mitra et al. (2019) where they found an expression for the death line
@@
To decide whether a pulsar is still active in radio, its critical spin down rate $\dot{P}_{\rm line}$ is computed knowing its spin period $P$. If the pulsar lies in the death valley defined by the inequality $10^{-0.55}<\dot{P}_{\rm line}/\dot{P}<10^{1.15}$ then a value for $\alpha_{l}$ is taken from a uniform distribution between $0\degr$ and $65\degr$. Moreover a value for $T_{6}$ is also taken from a uniform distribution between 1.9 and 2.8. The parameter $b$ is taken from a uniform distribution between 30 and 60. The boundary parameter values for the death valley were chosen to traverse the $P-\dot{P}$ diagram from the region in the bottom right quadrant where the density starts to decrease, to the detectable pulsar with the lowest $\dot{P}$ and highest $P$ in the same quadrant, as shown in Fig. 1. The edges of the death valley given by the values $10^{-0.55}$ and $10^{1.15}$ were found by taking the extreme values for the parameters that are drawn from the different uniform distributions. Then we compare its actual spin down rate $\dot{P}$, computed from the evolution model, to the $\dot{P}^{\rm new}_{\rm line}$ computed with equation (21) with the parameters which were drawn from uniform distributions. For pulsars that are not in the death valley, their $\dot{P}$ is only compared with the spin down rate $\dot{P}_{\rm line}$ obtained with the default parameters. If the pulsar has its $\dot{P}$, which was computed from the evolution model, below the one computed with the death line ($\dot{P}_{\rm line}$ or $\dot{P}^{\rm new}_{\rm line}$), then the pulsar is considered as dead, otherwise it is considered as still emitting photons. These choices were made in order to have more like a death valley near the death line than a strict condition of death, because the parameters $T_{6}$, $\alpha_{l}$ and $b$ may be different from one pulsar to another.
@@ 2.5 Detection
For each generated pulsar, we check whether it fulfills the criteria for detection depending on three factors. First, the beaming fraction indicates the fraction of the sky covered by the radiation beam and depends on the considered wavelength, here radio or $\gamma$-ray. The beaming fraction varies also with the pulsar spin rate, geometry and location of the emission regions. Before explaining how the beaming fraction in radio and in $\gamma$-ray are computed, it is important to define several angles.
@@
We choose a isotropic distribution for the Earth viewing angle $\xi$ as well as for the orientation of the unitary rotation vector. The Cartesian coordinates of the unit rotation vector ${n_{\Omega}}$ are $(\sin\theta_{n_{\Omega}}\cos\phi_{n_{\Omega}},\sin\theta_{n_{\Omega}}\sin\phi_{n_{\Omega}},\cos\theta_{n_{\Omega}})$. We set the Sun position at $(x_{\odot},y_{\odot},z_{\odot})=$ (0 kpc, 8.5 kpc, 15 pc) (Siegert, 2019). The coordinates for ${n}$ are
@@ 2.5.1 Radio detection model
The beaming fraction in radio depends on the half opening angle of the radio emission cone $\rho$ computed according to
@@
where $h_{\rm em}$ is the emission height, $P$ is the spin period of the pulsar and $c$ is the speed of light. The emission height is taken constant with an average value of $h_{\rm em}=3.10^{5}$ m, estimated from observations of a sample of pulsars by (Weltevrede & Johnston, 2008; Mitra, 2017; Johnston & Karastergiou, 2019; Johnston et al., 2023). The cone half-opening angle $\rho$ estimated in eq. (24) holds only for the last open field lines of a magnetic dipole and can only be applied for slow pulsars where the radio emission altitude is high enough for the multipolar components to decrease significantly and become negligible.
@@
The second factor for detection is the luminosity, which is also different between radio and gamma-rays. First, concerning the radio flux density, the formula used is the same as in Johnston et al. (2020) for a pulsar at 1.4 GHz, in order to model the detection carried out by the Parkes radio telescope in the Southern Galactic plane (Kramer et al., 2003; Lorimer et al., 2006; Cameron et al., 2020) and the Arecibo telescope in the northern plane (Cordes et al., 2006)
@@
$\tilde{w}_{r}=\sqrt{\left(w_{r}\frac{P}{2\pi}\right)^{2}+\tau_{samp}^{2}+\tau_{DM}^{2}+\tau_{scat}^{2}}$ (28a)
@@
Both $\tau_{DM}$ and $\tau_{scat}$ depend on the dispersion measure, which is found for each pulsar by running the code of Yao et al. (2017) converting the distance of a pulsar into the dispersion measure thanks to their state-of-the art model of the Galactic electron density distribution. The radio survey parameters used are the ones from the Parkes Multibeam Pulsar Survey (PMPS), see Table 4.
@@
2.5.2 Gamma-ray detection model
The $\gamma$-ray emission model relies on the striped wind model, describing gamma-ray photons production emanating from the current sheet within the striped wind. In order to detect the gamma-ray, the line of sight of the observer must remain around the equator plane with an inclination angle constrained by $|\xi-\pi/2|\leq\alpha$. The gamma-ray luminosity is extracted from a study of Kalapotharakos et al. (2019), where they showed that the luminosity is described by a fundamental plane. The 3D model depends on the magnetic field $B$, the spin down luminosity $\dot{E}$ and the cut-off energy $\epsilon_{cut}$. However, in our PPS the cut-off energy is not computed therefore we use their two dimensional version
@@ 2.5.2 Gamma-ray detection model
where $f_{\Omega}$ is a factor depending on the emission model reflecting the anisotropy. For the striped wind model Pétri (2011) showed that this factor varies between 0.22 and 1.90. Nevertheless, an approximation is made : if $\alpha<-\xi+0.6109$ then $f_{\Omega}=1.9$ otherwise $f_{\Omega}=1$. The pulsar is detected in gamma depending on the instrumental sensitivity as described below.
The sensitivity in gamma-ray is based on the expectation of the Fermi/LAT instrument  [^4] https://fermi.gsfc.nasa.gov/ssc/data/analysis/documentation/Cicerone
@@ 3 Results
Let us now move to the PPS simulation results. First we show runs simulating 10 million pulsars with and without a death line, allowing to simulate old pulsars (with the oldest aged $4.1\times 10^{8}$ yr). Then we show runs simulating 1 million pulsars (with the oldest aged $4.1\times 10^{7}$ yr) with and without considering a death line. We explore a possible spin-velocity misalignment effect due to the galactic potential and the impact of the initial spin period distribution, normal vs log-normal. Finally, the gamma-ray simulations and detection are discussed in details.
@@ 3.1.1 Without a death line
The first simulations performed with the Galactic potential implemented discarded the death line. The parameters used for these runs are given in Table 5. The total number of pulsars is summarized in Table 6.
@@
$\tau_{\text{birth}}$ (1/yr) $P_{\text{mean}}$ (ms) $B_{\text{mean}}$ (T) $\sigma_{p}$ $\sigma_{b}$ $\alpha_{d}$ $k_{\tau_{d}}$
41 129 $2.75\times 10^{8}$ 0.45 0.5 1.5 5
@@
¿31 11 0 8 3
¿28 161 9 75 77
Total 2653 2336 139 178
The quantities $N_{r}$, $N_{g}$, and $N_{rg}$ are the number of radio-only, gamma-only, and radio-loud gamma-ray pulsars, respectively, obtained from our simulation.
Compared to the results obtained by Dirson et al. (2022), at very high spin down luminosity, $\dot{E}>10^{31}$ W, the detection is almost identical but at $\dot{E}>10^{28}$ W there are much less pulsars detected in radio and radio/gamma in this work. There are probably two causes for this discrepancy in the number of detected pulsars at $\dot{E}>10^{28}$ W : firstly the difference in the spin period distribution at birth. Pulsars have a higher probability to start with a higher $P_{0}$ since the mean of the distribution is higher (in the previous study the mean was 60 ms while it is 129 ms now), therefore the slowing down of the pulsars can only decrease their rotation frequency $\Omega$, and equation (5) shows that $\dot{E}$ depends on $\Omega$. However, it does not affect the detection in gamma-rays because usually canonical gamma pulsars have a high $\dot{P}$ and a low $P$, being young, therefore their $\dot{E}$ is higher than for radio pulsars. Secondly, the ISM dispersion and scattering reduces this number of radio pulsars detected (the results without the ISM effect are not shown but without it there is much more radio pulsars detected). Nonetheless, in terms of total number of detected radio pulsars, the Galactic potential allows to have an increase compared to the number found in Dirson et al. (2022), a result expected because of the attractive nature of the potential of the Galaxy. In its previous version the PPS moved the pulsars in random spatial directions at constant speed, as a result no pulsar was bound to the Milky Way in closed orbits this is especially true for old pulsars which could have time to leave the Galaxy.
Figure 2: Distribution of the observed period taken from the ATNF catalogue, along with the simulations without the implementation of the death line. In green, the period of the simulated pulsars with age greater than $10^{7.5}$
Fig. 2 shows the distribution of the spin period of the detected pulsars in the simulation, in red and in observations, in blue. However the comparison is not satisfactory because there is an excess of pulsars detected in the simulation. Furthermore, this excess is also present in Fig. 3 where the age of the detected pulsars from the observations and simulation are plotted.
In order to verify if there is a correlation between old pulsars and the ones with a high spin period, we plotted in Fig. 2, in green the pulsars with an age greater than $10^{7.5}$ yr, corresponding to the oldest pulsars of the simulation and it corresponds to the excess found in Fig. 2. Due to the Galactic potential, more pulsars remain bound to the Galaxy, especially older pulsars which were escaping the Galaxy more easily in the work of Dirson et al. (2022) because of the absence of potential. Fig. 4 shows that many pulsars are in what is commonly called the graveyard of pulsars at the bottom right of the $P-\dot{P}$ diagram.
@@
Given this excess of long period pulsars without a death line, we decided to implement one, in order to stick closer to reality.

### v2

@@ 2.2 Pulsar evolution
The pulsar initial period $P_{0}$, inclination angle $\alpha_{0}$ and magnetic field $B_{0}$ are evolved in time in a fully self-consistent way taking into account the spin-down losses and the internal magnetic field dissipation within the neutron star. The neutron star period is evolved according to the force-free magnetosphere model which takes into account the electric current and charge flowing inside the magnetosphere. The spin down luminosity $\dot{E}$ is
@@
where $I$ is the neutron star moment of inertia approximately equal to $I\approx 10^{38}$ kg.m2, $\Omega=2\pi/P$ the rotation frequency of the pulsar ($P$ is the spin period), $\dot{\Omega}$ its time derivative, $R=12$ km the typical radius of a neutron star as found by recent NICER observations (Riley et al., 2019; Bogdanov et al., 2019), $B$ the magnetic field at the magnetic equator, $\alpha$ the inclination angle, $\mu_{0}$ the vacuum permeability constant, $\mu_{0}=4\pi\times 10^{-7}$ H/m and $c$ the speed of light. The term $L_{\rm ffe}$ for the spin down luminosity was given by Spitkovsky (2006) and Pétri (2012). Equation (6) clearly shows the correlation between the obliquity, the magnetic field and the rotation frequency. Combining equation (6) and (7) leads to
@@
where quantities with subscript 0 indicate their initial value and those without a subscript their current value at present time.
@@
where $\alpha_{d}$ is a constant parameter controlling the rate of the magnetic field decay and $\tau_{d}$ the typical decay timescale, which depends on the initial magnetic field as demonstrated by magneto-thermal evolution models (Viganò et al., 2013). We took these results into account by defining the decay timescale $\tau_{1}$ for a magnetic field strength $B_{1}$ such that $\tau_{1}\,B_{1}^{\alpha_{d}}=\tau_{d}\,B_{0}^{\alpha_{d}}$. Thus, $\tau_{1}$ is another magnetic field decay time scale, corresponding to a certain magnetic field value $B_{1}$. Three different $\tau_{1}$ are considered in this work and are randomly chosen for each pulsar generated : $1.5\times 10^{5}$ yr or $3.5\times 10^{5}$ yr or $2.5\times 10^{6}$ yr with each of these timescales corresponds to a magnetic field value $B_{1}$ of $1\times 10^{8}$ T, $3\times 10^{8}$ T and $2\times 10^{9}$ T (these values of $\tau_{1}$ and $B_{1}$ are taken from Fig. 10 of the work of Viganò et al. (2013)). Then the probabilities to get these values of $\tau_{1}$ and $B_{1}$ to be adopted for one pulsar are 0.23, 0.46 and 0.31 respectively (these probabilities were chosen as best educated guesses after several run with different values). The value of $\alpha_{d}$ can be found in Table 5 of Sect. 3.
In line with the spin down luminosity, the inclination angle satisfies another evolution equation that after integration was found by Phillipov et al. (2014) for a spherically symmetric neutron star with a constant magnetic field. With our decaying prescription the inclination angle $\alpha$ is found by solving
@@
where $t$ is the time representing the age of the pulsar and $K={R^{6}}/{I\,c^{3}}$. The typical decay timescale for a mean magnetic field of $2.5\times 10^{8}$ T is $4.6\times 10^{5}$ yr in Viganò (2013), however in this study the typical decay timescale associated to this magnetic field is $3.5\times 10^{5}$ yr which is very close to the estimate of Viganò (2013) for the same field strength.
2.3 Description of the Galactic potential
Finally, we need to follow the particle motion within the Galactic potential. Each pulsar evolves in the gravitational potential $\Phi$ subject to an acceleration ${\ddot{x}}$ according to
@@ 2.3 Description of the Galactic potential
Equation (13) is integrated numerically thanks to a Position Extended Forest Ruth-Like (PEFRL) algorithm (Young, 2014), a fourth order integration scheme, presented in Appendix A with a convergence and accuracy study. We now discuss the galactic potential model used.
@@
The potential for the dark matter halo $\Phi_{h}$, according to the frequently used potential of Navarro et al. (1997), is expressed as
@@ 2.4 The death line
Several equations for the death line have been proposed in the literature (Chen & Ruderman, 1993; Zhang et al., 2000; Gil & Mitra, 2001; Mitra et al., 2019). The most suitable one in our study is that of Mitra et al. (2019) where they found an expression relating $P$ and $\dot{P}$ by
@@
We allow the different parameters for this death valley to be in the following ranges : $\alpha_{l}$ is drawn from a uniform distribution between $0\degr$ and $65\degr$, $T_{6}$ is drawn from a uniform distribution between 1.9 and 2.8 and $b$ is drawn from a uniform distribution between 30 and 60. The boundary parameter values, corresponding to the edges of the death valley, were chosen to traverse the $P-\dot{P}$ diagram from the region in the bottom right quadrant where the number of pulsars starts to decrease (because this is likely the region where the death of pulsars begins to be considered, since less pulsars are there), to the detectable pulsar with the lowest $\dot{P}$ and highest $P$ in the same quadrant, as shown in Fig. 1. To decide whether a pulsar is still active in radio or not we apply the following criterion: if the pulsar lies in the death valley then all the parameters mentioned above are taken from uniform distributions in the ranges of the death valley in order to compute $\dot{P}^{\rm new}_{\rm line}$ with equation (22), the critical spin down rate of the considered pulsar. Then we compare its actual spin down rate $\dot{P}$, computed from the evolution model, to the $\dot{P}^{\rm new}_{\rm line}$. For pulsars that are not in the death valley, their $\dot{P}$ is only compared with the spin down rate $\dot{P}_{\rm line}$ obtained with the default parameters. If the pulsar has its $\dot{P}$, which was computed from the evolution model, below the one computed with the death line, then the pulsar is considered as dead, otherwise it is considered as still emitting photons. These choices were made in order to have more like a death valley near the death line than a strict condition of death, because the parameters $T_{6}$, $\alpha_{l}$ and $b$ may be different from one pulsar to another.
@@ 2.5 Detection
For each pulsar, we check whether it fulfills the detection criteria depending on three factors. First, the beaming fraction indicates the fraction of the sky covered by the radiation beam and depends on the considered wavelength, here radio or $\gamma$-ray. The beaming fraction varies also with the pulsar spin rate, geometry and location of the emission regions. Before explaining how the beaming fraction in radio and in $\gamma$-ray are computed, it is important to define several angles.
@@
We choose an isotropic distribution for the Earth viewing angle $\xi$ as well as for the orientation of the unitary rotation vector. The Cartesian coordinates of the unit rotation vector ${n_{\Omega}}$ are $(\sin\theta_{n_{\Omega}}\cos\phi_{n_{\Omega}},\sin\theta_{n_{\Omega}}\sin\phi_{n_{\Omega}},\cos\theta_{n_{\Omega}})$. We set the Sun position at $(x_{\odot},y_{\odot},z_{\odot})=$ (0 kpc, 8.5 kpc, 15 pc) (Siegert, 2019). The coordinates for ${n}$ are
@@ 2.5.1 Radio detection model
The beaming fraction in radio depends on the half opening angle of the radio emission cone $\rho$ computed according to Lorimer & Kramer (2004) by
@@
where $h_{\rm em}$ is the emission height, $P$ is the spin period of the pulsar and $c$ is the speed of light. The emission height is taken constant with an average value of $h_{\rm em}=3.10^{5}$ m, estimated from observations of a sample of pulsars by Weltevrede & Johnston (2008); Mitra (2017); Johnston & Karastergiou (2019); Johnston et al. (2023). The cone half-opening angle $\rho$ estimated in eq. (25) holds only for the last open field lines of a magnetic dipole and can only be applied for slow pulsars where the radio emission altitude is high enough for the multipolar components to decrease significantly and become negligible.
@@
The second factor for detection is the luminosity, which is also different between radio and $\gamma$-rays. First, concerning the radio flux density, the formula used is the same as in Johnston et al. (2020) for a pulsar at 1.4 GHz, in order to model the detection carried out by the Parkes radio telescope in the southern Galactic plane (Kramer et al., 2003; Lorimer et al., 2006; Cameron et al., 2020) and the Arecibo telescope in the northern plane (Cordes et al., 2006) (note that these two telescopes do not cover all the sky in the southern and northern plane, but we did not implemented any filter in the code)
@@
$\tilde{w}_{r}=\sqrt{\left({w_{r}\,P}/{2\pi}\right)^{2}+\tau_{samp}^{2}+\tau_{DM}^{2}+\tau_{scat}^{2}}$ (29a)
@@
Both $\tau_{DM}$ and $\tau_{scat}$ depend on the dispersion measure, which is found for each pulsar by running the code of Yao et al. (2017) converting the distance of a pulsar into the dispersion measure thanks to their state-of-the art model of the Galactic electron density distribution. The radio survey parameters used are the ones from the Parkes Multibeam Pulsar Survey (PMPS) taken from Manchester et al. (2001), see Table 4.
@@
2.5.2 $\gamma$-ray detection model
The $\gamma$-ray emission model relies on the striped wind model, describing $\gamma$-ray photons production emanating from the current sheet within the striped wind. In order to detect the $\gamma$-ray, the line of sight of the observer must remain around the equator plane with an inclination angle constrained by $|\xi-\pi/2|\leq\alpha$. The $\gamma$-ray luminosity is extracted from a study of Kalapotharakos et al. (2019), where they showed that the luminosity is described by a fundamental plane. The 3D model depends on the magnetic field $B$, the spin down luminosity $\dot{E}$ and the cut-off energy $\epsilon_{cut}$. However, in our PPS the cut-off energy is not computed therefore we use their two dimensional version
@@ 2.5.2 $\gamma$-ray detection model
where $f_{\Omega}$ is a factor depending on the emission model reflecting the anisotropy. For the striped wind model Pétri (2011) showed that this factor varies between 0.22 and 1.90. Nevertheless, an approximation is made : if $\alpha<-\xi+0.6109$ then $f_{\Omega}=1.9$ otherwise $f_{\Omega}=1$, since as can be observed in Fig. 7 of Pétri (2011) $f_{\Omega}$ is usually equal to one of this two values depending on the value of the inclination angle $\alpha$, hence this approximation. The pulsar is detected in gamma depending on the instrumental sensitivity as described below.
The sensitivity in $\gamma$-ray is based on the expectation of the Fermi/LAT instrument  [^5] https://fermi.gsfc.nasa.gov/ssc/data/analysis/documentation/Cicerone
@@ 3 Results
Let us now move to the PPS simulation results. First we show runs simulating 10 million pulsars with and without a death line, allowing to generate old pulsars with ages up to $4.1\times 10^{8}$ yr. Then we show runs simulating only 1 million pulsars with ages up to $4.1\times 10^{7}$ yr, also with and without a death line. We explore a possible spin-velocity misalignment effect due to the galactic potential and the impact of the initial spin period distribution, comparing normal vs log-normal distributions. Finally, the $\gamma$-ray simulations and detection are exposed in details.
@@ 3.1 Simulating 10 millions pulsars
Let us start with a sample containing 10 millions pulsars showing the impact of the death line. A same study will be done for 1 million pulsars.
@@ 3.1.1 Without a death line
The first set of simulations discarded the death line. The parameters used for these runs are given in Table 5. The total number of pulsars is summarized in Table 6.
@@
$\tau_{\text{birth}}$ (1/yr) $P_{\text{mean}}$ (ms) $B_{\text{mean}}$ (T) $\sigma_{p}$ $\sigma_{b}$ $\alpha_{d}$
41 129 $2.75\times 10^{8}$ 0.45 0.5 1.5
@@
¿31 7 0 4 3
¿28 161 4 87 63
Total 2491 2169 163 159
The quantities $N_{r}$, $N_{g}$, and $N_{rg}$ are the number of radio-only, $\gamma$-only, and radio-loud $\gamma$-ray pulsars, respectively, extracted from our simulation.
Compared to the results obtained by Dirson et al. (2022), at very high spin down luminosity, $\dot{E}>10^{31}$ W, the detection rate is almost identical but at $\dot{E}>10^{28}$ W much less pulsars are detected in radio and radio/$\gamma$-ray in our work. There are probably two causes for this discrepancy in the number of detected pulsars at high spin down luminosities $\dot{E}>10^{28}$ W : firstly the difference in the spin period distribution at birth. Pulsars have a higher probability to start with a higher $P_{0}$ since the mean of the distribution is higher (in the previous study the mean period was 60 ms while we set it to 129 ms), therefore the slowing down of the pulsars can only decrease their rotation frequency $\Omega$, and equation (6) shows that $\dot{E}$ depends on $\Omega$. However, it does not affect the detection rate in $\gamma$-rays because usually canonical $\gamma$-ray pulsars possess a high $\dot{P}$ and a low $P$, being young, therefore their $\dot{E}$ is higher than for radio pulsars. Secondly, the ISM dispersion and scattering reduces the number of detected radio pulsars, results without the ISM effect are not shown but discarding its impact would lead to much more detected radio pulsars. Nonetheless, in terms of total number of detected radio pulsars, including the Galactic potential increases this number compared to Dirson et al. (2022), a result expected because of the attractive nature of the potential of the Galaxy. Moreover in this latter version, the PPS moved pulsars in random spatial directions at constant speed, as a result no pulsar was bound to the Milky Way in closed orbits an artifact that is especially true for old pulsars which could have time to leave the Galaxy.
Figure 2: Distribution of the observed period taken from the ATNF catalogue, along with the simulations without the implementation of the death line. In green, the period of the simulated pulsars with characteristic age greater than $10^{8}$ yr.
Fig. 2 compares the distribution of simulated and observed spin periods, in red and blue respectively. However the comparison is not satisfactory because of an excess of pulsars detected in the simulation. Furthermore, this excess is also present in Fig. 3 where the characteristic age of the observed pulsars and the simulated pulsars are plotted. Note that we employ the characteristic ages defined by
$\tau_{c}=\frac{P}{2\dot{P}}$ (34)
because we usually do not know the real ages of the observed pulsars.
@@
In order to verify whether there is a correlation between old pulsars and those with a high spin period or not, Fig. 2 shows in green pulsars with a characteristic age greater than $10^{8}$ yr, corresponding to the oldest pulsars of the simulation. It coincides with the excess found in Fig. 2. We conclude that canonical pulsars older than $10^{8}$ yr measured as their real age in our simulations should be discarded at the end of the simulation, since the characteristic age is overestimating the real age of pulsars. Indeed canonical pulsars older than this do not exist. The Galactic potential forbids many pulsars to escape from the Milky Way, they remain bound to the Galaxy, in contrast to the older pulsars found in Dirson et al. (2022), that were escaping the Galaxy easily because of the absence of this potential. Fig. 4 shows that many pulsars are in what is commonly called the graveyard of pulsars at the bottom right of the $P-\dot{P}$ diagram, the death valley being shown in Fig. 4 but not implemented.
@@
Given this excess of long period pulsars when neglecting the death line, we decided to implement it, in order to stick closer to reality.